# Universal Solvability for Robot Motion Planning on Graphs

### v1

@@ 5 Deterministic Algorithm for USolR
In this section, we present a polynomial-time deterministic algorithm for the USolR problem, which is a derandomization of Algorithm 2 presented in Section 4.2. The algorithm uses as a subroutine, Algorithm $\mathcal{A}$ (discussed in Proposition 2.4) to solve intermediate instances of FRMP.
For a given USolR instance $(G(V,E)\text{, }p)$ to be universally solvable (or be a YES-instance), it must hold true that for any pair of configurations $(S,T),S,T\in S_{G}$, $T$ is reachable from $S$, i.e., $(S,T)\in\tilde{R}$. Thus, by Section 3, $Y_{S}$ and $Y_{T}$ (as defined in Definition 3.5) must also be mutually reachable. Hence, it is sufficient to examine the mutual reachability of configurations of the form $Y_{S}$ (for some configuration $S$) mapping the set $R$ of $p$ robots to $V_{p}$. There are $p!$ such configurations of the $p$ robots in $R$, each corresponding to a permutation $\pi\in\mathfrak{S}_{p}$ (the set of all permutation from $[p]$ to $[p]$), such that $S_{\pi}=S_{I}\circ\pi$ (i.e., $S_{\pi}(i)=v_{\pi(i)}$ for all $i\in[p]$). Thus, for a given instance to be universally solvable, all these configurations $\{S_{\pi}:R\to V_{p}\mid\pi\in\mathfrak{S}_{p}\}$ must be reachable from one another, or equivalently, all these configurations must be reachable from the identity configuration $S_{I}$, which corresponds to the identity permutation $\pi_{I}:[p]\to[p]$ defined as $\pi_{I}(i)=i$, for all $i\in[p]$.
@@
A naive approach to decide universal solvability could be invoking Algorithm $\mathcal{A}$ for the FRMP instance $(G(V,E)\text{, }p\text{, }(S_{I},S_{\pi}))$ corresponding to each of these $p!$ configurations (checking the reachability of each $S_{\pi}$ from the configuration $S_{I}$). However, this approach would have time complexity $\mathcal{O}(p!\cdot(|V|+|E|))$, which is super-polynomial in the size of the input parameter $p$. However, in the following lemma, we now argue that it is sufficient to check the reachability from $S_{I}$ of $p-1$ specific configurations from the set $\{S_{\pi}:R\to V_{p}\mid\pi\in\mathfrak{S}_{p}\}$. Consider the set of configurations $\{S_{\pi^{*}_{t}}\mid t\in[p-1]\}$, where the permutation $\pi^{*}_{t}$ is defined as follows.
@@
[…]
Now, we repeat the same procedure for the robot that occupies the vertex $v_{p-1}$ in $S_{\pi}$. Let this robot be denoted as $j^{\prime}=\pi^{-1}(v_{p-1})$. Then, in the current configuration $S_{1}$, robot $j^{\prime}$ occupies the vertex $S_{1}(j^{\prime})$. Now, we can again move the robot from $S_{1}(j^{\prime})$ to $v_{p-1}$ following the same procedure as above. We can keep repeating this process, and reach configurations $S_{2},S_{3},..,S_{p}(=S_{\pi})$ sequentially using valid moves. Hence, we can reach the configuration $S_{\pi}$ from $S_{I}$ for any arbitrary $\pi$. This shows that $S_{I}\sim S_{\pi}$, for all permutations $\pi:[p]\to[p]$. From Section 5, this implies $G$ is universally solvable for $p$ robots. This completes the proof of the backward direction, thereby proving the lemma.
With this background, we introduce our deterministic algorithm, which we refer to as Algorithm 3 and which derandomizes Algorithm 2. We state the following theorem.
@@ 6 Optimized Deterministic Algorithm for USolR
In the previous section, we designed and analysed a polynomial-time deterministic algorithm solving USolR for $p$ robots and an input graph $G(V,E)$, with a running time guarantee of $\mathcal{O}(p\cdot(|V|+|E|))$. Compared to the randomized algorithm, the running time blows up by a factor of $p$. In this section, we provide an optimised version of Algorithm 3, that exploits the structural properties of the input graph.
We start of by recalling Theorem 5 in the paper due to Yu and Rus [24]. We state this as the following proposition, as per the notation we introduced in this paper.
@@
Let $S:R\to V$ and $T:R\to V$ be arbitrary configurations. Due to Lemma 3.6, there exists a configuration $S^{\prime}:R\to C^{\prime}$ such that $S\sim S^{\prime}$. Similarly, there exists a configuration $T^{\prime}:R\to C^{\prime}$ such that $T\sim T^{\prime}$. Further, due to Proposition 6.1, $S^{\prime}$ and $T^{\prime}$ are reachable from each other, just by using the vertices and edges in the graph induced by $C$; this gives $S^{\prime}\sim T^{\prime}$. Putting everything together, we have $S\sim S^{\prime}\sim T^{\prime}\sim T$. This holds true for any pair of arbitrary configurations $S$ and $T$, hence $G$ is universally solvable for $p$ robots.
@@
This coupled with Lemma 6.2 gives us the following observation.
@@
Let us consider the instance $G(V,E)$, $p$ of USolR. If $G$ is disconnected, it is not universally solvable. Otherwise, by Section 6, whenever $|E|\geq p\cdot|V|$, the input instance must be a universally solvable. Therefore, we can check if $|E|\geq p\cdot|V|$ and output YES if the check passes. Otherwise we run Algorithm 3 on the input instance to get the answer.
The correctness is of this algorithm is due to Section 6 and the correctness of Algorithm 3. All that remains is to analyse the running time of this algorithm.
If $G$ is disconnected, we detect in linear time, $\mathcal{O}(|V|+|E|)$, and output NO correctly. If $|E|\geq p\cdot|V|$, we take input the graph, perform the check, and output YES correctly. This takes linear time, $\mathcal{O}(|V|+|E|)$, in total. Otherwise, if $|E|<p\cdot|V|$, then we run Algorithm 3, which takes time $\mathcal{O}(p\cdot(|V|+|E|))$ (Theorem 5.3). This completes the proof.
The pseudocode of this algorithm, called Algorithm 4 can be found in the Appendix. Algorithm 4 performs at least as good as Algorithm 3 in all graphs. Moreover, when $p$ is small, and the graph is dense enough, Algorithm 4 utilises Section 6 to output YES correctly. Therefore, the speed up is significant for dense graphs with small number of robots.
@@ 7 Augmenting Graphs for Universal Solvability
For the problem to be well-defined, we specify the number of robots $p$ for which we want $G$ to be universally solvable as part of the input. For the warehouse setting, where we would ideally want universal solvability for any number of robots $p\leq|V|$, simply solving the above problem for $p=|V|$ guarantees universal solvability for all $p\leq|V|$. This is because, for any configuration $S$ with the number of robots $p<|V|$, we can consider a configuration $S^{\prime}$ of $|V|$ robots, with $p$ robots occupying the same vertices as in $S$ and $(|V|-p)$ dummy robots occupying the remaining $(|V|-p)$ empty vertices of $S$. Thus, if we can route between any two configurations of $|V|$ robots, we can route between any two configurations of a subset of size $p$ of $|V|$ robots, by transforming them into configurations of size $|V|$ by adding $(|V|-p)$ dummy robots.
We study the EAUS problem only for connected graphs and further assume that the graph $G$ has at least four vertices. For the case where the number of vertices $|V|$ is less than or equal to three, a simplistic brute force approach can be used to solve EAUS. On general graphs we do a case analysis for different classes of graphs based on vertex and edge connectivity. We rely upon certain structural results related to universal solvability discussed by Yu and Rus [24] for our analysis. We study combinatorial bounds on the number of edge additions $\beta$ for the problem for the different graph classes below.
@@
Previously, we discussed that the class of 2-connected graphs, that are not simple cycles, is universally solvable for all $p$ (Proposition 6.1). Therefore, these graphs do not require any edge additions for achieving universal solvability. The remaining case is that of simple cycles.
@@
For the case, when the graph $G(V,E)$ is a simple cycle, it is possible to make $G$ universally solvable for all $p\leq|V|$ with just one extra edge addition. We can add the extra edge as any chord of the cycle $G$, thus forming a 2-connected component of size $|V|$, which is not a cycle. By Proposition 6.1, this becomes universally solvable. Therefore, an edge budget $\beta=1$ suffices for achieving universal solvability of simple cycles.
@@
Any graph $G$ that is 2-edge connected but not 2-connected, is universally solvable for all $p<|V|$, as proved in Lemma 12 in Yu and Rus [24]. Thus, if such a graph is not universally solvable, it must be that $p=|V|$. We therefore discuss the EAUS problem for this class of graphs for $p=|V|$ only. Since $G$ is 2-edge connected but not 2-connected, there are no cut edges in $G$, and there is at least one cut vertex in $G$. Now, before discussing this graph class, we first state a lemma that we need for our analysis.
@@
Firstly, we discuss the case when all the 2-connected components are simple cycles. Then, if all the 2-connected components are odd-length cycles, then the graph $G$ is not universally solvable for $p=|V|$, while even if one of the cycles is an even-length cycle, then the graph $G$ is universally solvable for $p=|V|$. This is because in the case of 2-edge connected but not 2-connected graphs, if all 2-connected components are odd-length cycles, all configurations are partitioned into two equivalence classes which are unreachable for each other, such that two configurations which are identical except any two adjacent vertices swapped, lie in different equivalence classes. However, if at least one of the cycles is of even length, it allows two configurations which are identical except two vertices swapped, to become reachable from each other, thus collapsing the two equivalence classes into one class. The detailed proof of this has been discussed in Theorem 2 of [24].
Now, we consider the general case when at least one 2-connected component is not a simple cycle. Then, such a 2-connected component must contain at least 4 vertices. Now, consider one such 2-connected component $g$ of size at least 4 which is not a cycle, and denote its size as $s_{g}$. Then, we can route between any two configurations of $s_{g}$ robots in $g$ (Proposition 6.1). In particular, we can route between two configurations of $s_{g}$ robots, which are identical except that the robots on two adjacent vertices are swapped. Now, consider the configurations of $|V|$ robots over the entire graph $G$. Then, any two configurations of $|V|$ robots, which are identical except robots on two adjacent vertices of $g$ swapped, are also reachable from each other. Thus, the two equivalence classes of reachability (as there are in the case when all 2-connected components are odd-length cycles) again collapse into one class, implying universal solvability of $G$ for $p=|V|$.
@@
Now, we first consider the special case where all the 2-connected components in $G$ are just triangles or $3$-cycles. For $p=|V|$, this graph class is not universally solvable (due to Lemma 7.1). To make it universally solvable for $p=|V|$, we consider two adjacent triangles $T_{1}$ and $T_{2}$ that share a common vertex, and add an edge joining one vertex of $T_{1}$ (one of the two vertices that are not shared with $T_{2}$) with one vertex of $T_{2}$ (again, one of the two vertices that not shared with $T_{1}$). This creates a 2-connected component of size 5 which is not a cycle. By Lemma 7.1 this makes the graph universally solvable. Hence, adding just 1 extra edge (i.e., $\beta=1$) suffices to make this graph class solvable.
Thus, the other case that remains is when at least one of the 2-connected components of $G$ has size 4 or more. Now, for $G$ to be not universally solvable for $p=|V|$, all such 2-connected components of size 4 or more must be only simple odd-length cycles (due to Lemma 7.1). We can add one extra edge to any one of these cycles to create a 2-connected component of size at least 4 that is not a simple cycle, making $G$ universally solvable for $p$ robots (again due to Lemma 7.1).
@@
We first prove an upper bound on the edge addition budget $\beta$ to achieve universal solvability for $p$ robots. We show that for a given instance $(G(V,E),p)$ which is not universally solvable, where $G$ is a 1-edge connected graph, at most $p-2$ edge additions suffice to make $G$ universally solvable for $p$ robots (i.e., $\beta\leq p-2$). Clearly, since $G$ is not universally solvable for given $p$, $G$ must not have any 2-connected component which is not a cycle of size at least $p$ (Lemma 6.2). Now, consider any vertex $w$ in $G$ with degree at least 2. Let $x$ and $y$ be two of its neighbours. Consider $p-3$ other vertices and fix an arbitrary order of these $p-3$ vertices as $v_{4},v_{5},...,v_{p}$. Now consider the vertex sequence $(x(=v_{1}),w(=v_{2}),y(=v_{3}),v_{4},...,v_{p})$. $G$ already has $\{w,x\}$ and $\{w,y\}$ as edges. Now, we add all edges connecting adjacent vertices $\{v_{i},v_{i+1}\}$ (for convenience, define $v_{p+1}:=v_{1}$) for all $i\in\{3,4...,p\}$ to form a cycle. Some of these edges may already be present, so we add a maximum of $p-2$ edges this way. Now, we analyse the case of $p=|V|$ and $p<|V|$ separately.
  • For the case $p=|V|$, consider the structure of $G$ before we augmented $G$ to form the $p$-sized cycle. Since $|V|\geq 4$ and $G$ is connected, therefore, $G$ must have at least 3 edges (before addition of the extra edges). Two of these edges are $\{w,x\}$ and $\{w,y\}$. Therefore, there must have been at least one more edge in $G$ before the addition of extra edges. Now, if that edge connected two adjacent vertices $\{v_{i},v_{i+1}\}$ for some $i\in\{3,4,...,p\}$ (again, define $v_{p+1}:=v_{1}$), then we would have added at most $p-3$ edges to form the cycle. Then, we could add an extra edge between any two non-adjacent vertices of the cycle we created to form a chord, thus resulting in a 2-connected component of size $p$, which is not a cycle. Thus, we add a maximum of $p-2$ extra edges. If the extra edge was not between any two adjacent vertices of the cycle, then clearly, it must have been a chord between two non-adjacent vertices of the cycle (since all vertices are part of the cycle), and therefore, we again have a 2-connected component of size $p$ which is not a simple cycle. This makes the graph universally solvable (Proposition 6.1). Thus, we can achieve universal solvability for a general graph $G(V,E)$ and $p=|V|$ with a budget $\beta=p-2$ for edge additions in the worst case.
@@
Now, we provide an asymptotically matching lower bound of $\Theta(p)$, even for the restricted class of 1-edge connected graphs. Consider the case $p=|V|$. Consider the star graph $G(V,E)$ with $|V|-1$ leaves. Note that for $p=|V|$, a robot can move only if it is located on a cycle. Since there are $|V|-1$ leaves, for all robots in these leaves to even be able to move, we need to at least make each leaf, a part of some cycle. This would require at least $(|V|-1)/2=\Theta(|V|)=\Theta(p)$ additional edges. Moreover, for any tree with $\Theta(|V|)=\Theta(p)$ many leaves, at least $\Theta(|V|)=\Theta(p)$ additional edges would be required to make the graph universally solvable. We get a lower bound of $\beta\geq(|V|-1)/2$.

### v2

@@ 5 Deterministic Algorithm for USolR
In this section, we present a polynomial-time deterministic algorithm for the USolR problem, which is a derandomization of Algorithm˜2 presented in Section˜4.2. The algorithm uses as a subroutine, Algorithm $\mathcal{A}$ (discussed in Proposition˜2.4) to solve intermediate instances of FRMP.
For a given USolR instance $(G(V,E)\text{, }p)$ to be universally solvable (or be a YES-instance), it must hold true that for any pair of configurations $(S,T),S,T\in S_{G}$, $T$ is reachable from $S$, i.e., $(S,T)\in\tilde{R}$. Thus, by Section˜3, $Y_{S}$ and $Y_{T}$ (as defined in Definition˜3.5) must also be mutually reachable. Hence, it is sufficient to examine the mutual reachability of configurations of the form $Y_{S}$ (for some configuration $S$) mapping the set $R$ of $p$ robots to $V_{p}$. There are $p!$ such configurations of the $p$ robots in $R$, each corresponding to a permutation $\pi\in\mathfrak{S}_{p}$ (the set of all permutation from $[p]$ to $[p]$), such that $S_{\pi}=S_{I}\circ\pi$ (i.e., $S_{\pi}(i)=v_{\pi(i)}$ for all $i\in[p]$). Thus, for a given instance to be universally solvable, all these configurations $\{S_{\pi}:R\to V_{p}\mid\pi\in\mathfrak{S}_{p}\}$ must be reachable from one another, or equivalently, all these configurations must be reachable from the identity configuration $S_{I}$, which corresponds to the identity permutation $\pi_{I}:[p]\to[p]$ defined as $\pi_{I}(i)=i$, for all $i\in[p]$.
@@
A naive approach to decide universal solvability could be invoking Algorithm $\mathcal{A}$ for the FRMP instance $(G(V,E)\text{, }p\text{, }(S_{I},S_{\pi}))$ corresponding to each of these $p!$ configurations (checking the reachability of each $S_{\pi}$ from the configuration $S_{I}$). However, this approach would have time complexity $\mathcal{O}(p!\cdot(|V|+|E|))$, which is super-polynomial in the size of the input parameter $p$. However, in the following lemma, we now argue that it is sufficient to check the reachability from $S_{I}$ of p1p-1 specific configurations from the set $\{S_{\pi}:R\to V_{p}\mid\pi\in\mathfrak{S}_{p}\}$. Consider the set of configurations $\{S_{\pi^{*}_{t}}\mid t\in[p-1]\}$, where the permutation $\pi^{*}_{t}$ is defined as follows.
@@
For the backward direction, we are given that $S_{I}\sim S_{\pi^{*}_{t}}$ for all $t\in[p-1]$. This implies that $S_{\pi}(=S_{I}\circ\pi)\sim S_{\pi^{*}_{t}}\circ\pi$, for all $t\in[p-1]$ for any permutation $\pi$ (Section˜2.1). This signifies that for any configuration $S_{\pi}$, using a sequence of valid moves, we can swap the robots occupying vertices $v_{t}$ and $v_{t+1}$ (while keeping other robots fixed), for every $t\in[p-1]$. To prove that an instance is universally solvable, we need to show that $S_{I}\sim S_{\pi}$ for all $\pi\in\mathfrak{S}_{p}$ (Section˜5). Now, we show that $S_{I}\sim S_{\pi}$ for any $\pi\in\mathfrak{S}_{p}$ given $S_{I}\sim S_{\pi^{*}_{t}}$ for all $t\in[p-1]$. Consider any arbitrary permutation $\pi\in\mathfrak{S}_{p}$ and the corresponding configuration $S_{\pi}$. We incrementally reach the configuration $S_{\pi}$ from $S_{I}$ by routing the robots in the following manner: we first move the robot in $S_{I}$ that occupies the vertex $v_{p}$ in $S_{\pi}$. Let this robot be denoted as $j=\pi^{-1}(v_{p})$. Then, in $S_{I}$, robot $j$ occupies $v_{j}$. Therefore, we first move robot $j$ from $v_{j}$ to $v_{p}$. We argued that it is possible to swap robots in $v_{t}$ and $v_{t+1}$ for all $t\in[p-1]$ using a sequence of valid moves. We use this to establish that we can route the robot $j$ from $v_{j}$ to $v_{p}$. For that, initially, we have the configuration $S_{0}=S_{I}$. The vertices $v_{j},v_{j+1},...,v_{p}$ are occupied by robots j,j+1,,pj,j+1,...,p respectively. We can reach the configuration $S_{\pi^{*}_{j}}$ from $S_{I}$ (i.e., only the robots at vertices $v_{j}$ and $v_{j+1}$ are exchanged). Now, we have robots arranged according to the configuration $S_{\pi^{*}_{j}}$. Further, we can exchange the robots at vertices $j+1$ and $j+2$ in this configuration as $S_{\pi^{*}_{j}}(=S_{I}\circ\pi^{*}_{j})\sim S_{\pi^{*}_{j+1}}\circ\pi^{*}_{j}$ as discussed above. This can be repeated for j+2,j+3,..,p1j+2,j+3,..,p-1, so that now robot $j$ is at vertex $v_{p}$. In the current configuration $S_{1}$ (obtained by placing $1$ robot at its correct position in $S_{\pi}$), the robots 1,2,..,j1,j,j+1,..,p1,2,..,j-1,j,j+1,..,p occupy vertices $v_{1},v_{2},..,v_{j-1},v_{p},v_{j},..,v_{p-1}$ in the specific order.
Now, we repeat the same procedure for the robot that occupies the vertex $v_{p-1}$ in $S_{\pi}$. Let this robot be denoted as $j^{\prime}=\pi^{-1}(v_{p-1})$. Then, in the current configuration $S_{1}$, robot $j^{\prime}$ occupies the vertex $S_{1}(j^{\prime})$. Now, we can again move the robot from $S_{1}(j^{\prime})$ to $v_{p-1}$ following the same procedure as above. We can keep repeating this process, and reach configurations $S_{2},S_{3},..,S_{p}(=S_{\pi})$ sequentially using valid moves. Hence, we can reach the configuration $S_{\pi}$ from $S_{I}$ for any arbitrary $\pi$. This shows that $S_{I}\sim S_{\pi}$, for all permutations $\pi:[p]\to[p]$. From Section˜5, this implies $G$ is universally solvable for $p$ robots. This completes the proof of the backward direction, thereby proving the lemma.
With this background, we introduce our deterministic algorithm, which we refer to as Algorithm˜3 and which derandomizes Algorithm˜2. We state the following theorem.
@@ 6 Optimized Deterministic Algorithm for USolR
In the previous section, we designed and analysed a polynomial-time deterministic algorithm solving USolR for $p$ robots and an input graph $G(V,E)$, with a running time guarantee of $\mathcal{O}(p\cdot(|V|+|E|))$. Compared to the randomized algorithm, the running time blows up by a factor of $p$. In this section, we provide an optimised version of Algorithm˜3, that exploits the structural properties of the input graph.
We start of by recalling Theorem 5 in the paper due to Yu and Rus [yu2014pebblemotiongraphsrotations]. We state this as the following proposition, as per the notation we introduced in this paper.
@@
Let $S:R\to V$ and $T:R\to V$ be arbitrary configurations. Due to Lemma˜3.6, there exists a configuration $S^{\prime}:R\to C^{\prime}$ such that $S\sim S^{\prime}$. Similarly, there exists a configuration $T^{\prime}:R\to C^{\prime}$ such that $T\sim T^{\prime}$. Further, due to Proposition˜6.1, $S^{\prime}$ and $T^{\prime}$ are reachable from each other, just by using the vertices and edges in the graph induced by $C$; this gives $S^{\prime}\sim T^{\prime}$. Putting everything together, we have $S\sim S^{\prime}\sim T^{\prime}\sim T$. This holds true for any pair of arbitrary configurations $S$ and $T$, hence $G$ is universally solvable for $p$ robots.
@@
This coupled with Lemma˜6.2 gives us the following observation.
@@
Let us consider the instance $G(V,E)$, $p$ of USolR. If $G$ is disconnected, it is not universally solvable. Otherwise, by Section˜6, whenever $|E|\geq p\cdot|V|$, the input instance must be a universally solvable. Therefore, we can check if $|E|\geq p\cdot|V|$ and output YES if the check passes. Otherwise we run Algorithm˜3 on the input instance to get the answer.
The correctness is of this algorithm is due to Section˜6 and the correctness of Algorithm˜3. All that remains is to analyse the running time of this algorithm.
If $G$ is disconnected, we detect in linear time, $\mathcal{O}(|V|+|E|)$, and output NO correctly. If $|E|\geq p\cdot|V|$, we take input the graph, perform the check, and output YES correctly. This takes linear time, $\mathcal{O}(|V|+|E|)$, in total. Otherwise, if $|E|<p\cdot|V|$, then we run Algorithm˜3, which takes time $\mathcal{O}(p\cdot(|V|+|E|))$ (Theorem˜5.3). This completes the proof.
The pseudocode of this algorithm, called Algorithm˜4 can be found in the Appendix. Algorithm˜4 performs at least as good as Algorithm˜3 in all graphs. Moreover, when $p$ is small, and the graph is dense enough, Algorithm˜4 utilises Section˜6 to output YES correctly. Therefore, the speed up is significant for dense graphs with small number of robots.
@@ 7 Augmenting Graphs for Universal Solvability
For the problem to be well-defined, we specify the number of robots $p$ for which we want $G$ to be universally solvable as part of the input. For the warehouse setting, where we would ideally want universal solvability for any number of robots $p\leq|V|$, simply solving the above problem for $p=|V|$ guarantees universal solvability for all $p\leq|V|$. This is because, for any configuration $S$ with the number of robots $p<|V|$, we can consider a configuration $S^{\prime}$ of $|V|$ robots, with $p$ robots occupying the same vertices as in $S$ and (|V|p)(|V|-p) dummy robots occupying the remaining (|V|p)(|V|-p) empty vertices of $S$. Thus, if we can route between any two configurations of $|V|$ robots, we can route between any two configurations of a subset of size $p$ of $|V|$ robots, by transforming them into configurations of size $|V|$ by adding (|V|p)(|V|-p) dummy robots.
We study the EAUS problem only for connected graphs and further assume that the graph $G$ has at least four vertices. For the case where the number of vertices $|V|$ is less than or equal to three, a simplistic brute force approach can be used to solve EAUS. On general graphs we do a case analysis for different classes of graphs based on vertex and edge connectivity. We rely upon certain structural results related to universal solvability discussed by Yu and Rus [yu2014pebblemotiongraphsrotations] for our analysis. We study combinatorial bounds on the number of edge additions $\beta$ for the problem for the different graph classes below.
@@
Previously, we discussed that the class of 2-connected graphs, that are not simple cycles, is universally solvable for all $p$ (Proposition˜6.1). Therefore, these graphs do not require any edge additions for achieving universal solvability. The remaining case is that of simple cycles.
@@
For the case, when the graph $G(V,E)$ is a simple cycle, it is possible to make $G$ universally solvable for all $p\leq|V|$ with just one extra edge addition. We can add the extra edge as any chord of the cycle $G$, thus forming a 2-connected component of size $|V|$, which is not a cycle. By Proposition˜6.1, this becomes universally solvable. Therefore, an edge budget $\beta=1$ suffices for achieving universal solvability of simple cycles.
@@
Any graph $G$ that is 2-edge connected but not 2-connected, is universally solvable for all $p<|V|$, as proved in Lemma 12 in Yu and Rus [yu2014pebblemotiongraphsrotations]. Thus, if such a graph is not universally solvable, it must be that $p=|V|$. We therefore discuss the EAUS problem for this class of graphs for $p=|V|$ only. Since $G$ is 2-edge connected but not 2-connected, there are no cut edges in $G$, and there is at least one cut vertex in $G$. Now, before discussing this graph class, we first state a lemma that we need for our analysis.
@@
Firstly, we discuss the case when all the 2-connected components are simple cycles. Then, if all the 2-connected components are odd-length cycles, then the graph $G$ is not universally solvable for $p=|V|$, while even if one of the cycles is an even-length cycle, then the graph $G$ is universally solvable for $p=|V|$. This is because in the case of 2-edge connected but not 2-connected graphs, if all 2-connected components are odd-length cycles, all configurations are partitioned into two equivalence classes which are unreachable for each other, such that two configurations which are identical except any two adjacent vertices swapped, lie in different equivalence classes. However, if at least one of the cycles is of even length, it allows two configurations which are identical except two vertices swapped, to become reachable from each other, thus collapsing the two equivalence classes into one class. The detailed proof of this has been discussed in Theorem 2 of [yu2014pebblemotiongraphsrotations].
Now, we consider the general case when at least one 2-connected component is not a simple cycle. Then, such a 2-connected component must contain at least 4 vertices. Now, consider one such 2-connected component $g$ of size at least 4 which is not a cycle, and denote its size as $s_{g}$. Then, we can route between any two configurations of $s_{g}$ robots in $g$ (Proposition˜6.1). In particular, we can route between two configurations of $s_{g}$ robots, which are identical except that the robots on two adjacent vertices are swapped. Now, consider the configurations of $|V|$ robots over the entire graph $G$. Then, any two configurations of $|V|$ robots, which are identical except robots on two adjacent vertices of $g$ swapped, are also reachable from each other. Thus, the two equivalence classes of reachability (as there are in the case when all 2-connected components are odd-length cycles) again collapse into one class, implying universal solvability of $G$ for $p=|V|$.
@@
Now, we first consider the special case where all the 2-connected components in $G$ are just triangles or $3$-cycles. For $p=|V|$, this graph class is not universally solvable (due to Lemma˜7.1). To make it universally solvable for $p=|V|$, we consider two adjacent triangles $T_{1}$ and $T_{2}$ that share a common vertex, and add an edge joining one vertex of $T_{1}$ (one of the two vertices that are not shared with $T_{2}$) with one vertex of $T_{2}$ (again, one of the two vertices that not shared with $T_{1}$). This creates a 2-connected component of size 5 which is not a cycle. By Lemma˜7.1 this makes the graph universally solvable. Hence, adding just 1 extra edge (i.e., $\beta=1$) suffices to make this graph class solvable.
Thus, the other case that remains is when at least one of the 2-connected components of $G$ has size 4 or more. Now, for $G$ to be not universally solvable for $p=|V|$, all such 2-connected components of size 4 or more must be only simple odd-length cycles (due to Lemma˜7.1). We can add one extra edge to any one of these cycles to create a 2-connected component of size at least 4 that is not a simple cycle, making $G$ universally solvable for $p$ robots (again due to Lemma˜7.1).
@@
We first prove an upper bound on the edge addition budget $\beta$ to achieve universal solvability for $p$ robots. We show that for a given instance $(G(V,E),p)$ which is not universally solvable, where $G$ is a 1-edge connected graph, at most p2p-2 edge additions suffice to make $G$ universally solvable for $p$ robots (i.e., $\beta\leq p-2$). Clearly, since $G$ is not universally solvable for given $p$, $G$ must not have any 2-connected component which is not a cycle of size at least $p$ (Lemma˜6.2). Now, consider any vertex $w$ in $G$ with degree at least 2. Let $x$ and $y$ be two of its neighbours. Consider p3p-3 other vertices and fix an arbitrary order of these p3p-3 vertices as $v_{4},v_{5},...,v_{p}$. Now consider the vertex sequence $(x(=v_{1}),w(=v_{2}),y(=v_{3}),v_{4},...,v_{p})$. $G$ already has $\{w,x\}$ and $\{w,y\}$ as edges. Now, we add all edges connecting adjacent vertices $\{v_{i},v_{i+1}\}$ (for convenience, define $v_{p+1}:=v_{1}$) for all $i\in\{3,4...,p\}$ to form a cycle. Some of these edges may already be present, so we add a maximum of p2p-2 edges this way. Now, we analyse the case of $p=|V|$ and $p<|V|$ separately.
  • For the case $p=|V|$, consider the structure of $G$ before we augmented $G$ to form the $p$-sized cycle. Since $|V|\geq 4$ and $G$ is connected, therefore, $G$ must have at least 3 edges (before addition of the extra edges). Two of these edges are $\{w,x\}$ and $\{w,y\}$. Therefore, there must have been at least one more edge in $G$ before the addition of extra edges. Now, if that edge connected two adjacent vertices $\{v_{i},v_{i+1}\}$ for some $i\in\{3,4,...,p\}$ (again, define $v_{p+1}:=v_{1}$), then we would have added at most p3p-3 edges to form the cycle. Then, we could add an extra edge between any two non-adjacent vertices of the cycle we created to form a chord, thus resulting in a 2-connected component of size $p$, which is not a cycle. Thus, we add a maximum of p2p-2 extra edges. If the extra edge was not between any two adjacent vertices of the cycle, then clearly, it must have been a chord between two non-adjacent vertices of the cycle (since all vertices are part of the cycle), and therefore, we again have a 2-connected component of size $p$ which is not a simple cycle. This makes the graph universally solvable (Proposition˜6.1). Thus, we can achieve universal solvability for a general graph $G(V,E)$ and $p=|V|$ with a budget $\beta=p-2$ for edge additions in the worst case.
@@
Now, we provide an asymptotically matching lower bound of $\Theta(p)$, even for the restricted class of 1-edge connected graphs. Consider the case $p=|V|$. Consider the star graph $G(V,E)$ with |V|1|V|-1 leaves. Note that for $p=|V|$, a robot can move only if it is located on a cycle. Since there are |V|1|V|-1 leaves, for all robots in these leaves to even be able to move, we need to at least make each leaf, a part of some cycle. This would require at least $(|V|-1)/2=\Theta(|V|)=\Theta(p)$ additional edges. Moreover, for any tree with $\Theta(|V|)=\Theta(p)$ many leaves, at least $\Theta(|V|)=\Theta(p)$ additional edges would be required to make the graph universally solvable. We get a lower bound of $\beta\geq(|V|-1)/2$.